Grade 11Algebra

if logb2 =a and log3b=c then find the value of logb500

Naman
9 Years agoGrade 11

1 Answer

Samyak Gupta
9 Years ago
We can write 500 as 2*2*5*5*5Then logb^500 = logb ^(2*2*5*5*5) But logax +logay= log a^xySo logb^500 = logb4+ logb125 = logb^2² + logb^5³ = 2a + 3c